OpenStudy (anonymous):

A quarterback for the Seattle Seahawks completes 54% of his passes. Let the random variable X be the number of passes completed in 20 attempts. a) What is the probability that the quarterback throws three passes before he gets his first completion (on his fourth attempt)? b) How many passes should the quarterback expect to make before he gets his first completion? c) What is the probability that he has to throw more than 5 passes to get his first completion?

OpenStudy (anonymous):

@jim_thompson5910

jimthompson5910 (jim_thompson5910):

this is another geometric distribution problem

jimthompson5910 (jim_thompson5910):

any ideas here?

OpenStudy (anonymous):

hmm not really, sorry

jimthompson5910 (jim_thompson5910):

we're using the same formula Pr(X = k) = (1-p)^(k-1)*p

jimthompson5910 (jim_thompson5910):

but now for this problem, the value of p is 0.54 so p = 0.54 for each part a) through c)

jimthompson5910 (jim_thompson5910):

in part a), the value of k is k = 4

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

how do i set it up

jimthompson5910 (jim_thompson5910):

Pr(X = k) = (1-p)^(k-1)*p Pr(X = k) = (1-0.54)^(k-1)*0.54 Pr(X = 4) = (1-0.54)^(4-1)*0.54 Pr(X = 4) = ???

OpenStudy (anonymous):

my calculator hasnt been wokring right on these...can u help out plz?

jimthompson5910 (jim_thompson5910):

use google as a calculator then http://www.google.com/search?hl=&q=%281-0.54%29^%284-1%29*0.54&sourceid=navclient-ff&rlz=1B3GGLL_enUS420US420&ie=UTF-8

OpenStudy (anonymous):

0.05256144

jimthompson5910 (jim_thompson5910):

yep

jimthompson5910 (jim_thompson5910):

that's the probability that the quarterback throws three passes before he gets his first completion (on his fourth attempt)

OpenStudy (anonymous):

kk! so, thats a? now b

jimthompson5910 (jim_thompson5910):

part b evaluate 1/p = 1/0.54 = ??? and round up

OpenStudy (anonymous):

2

OpenStudy (anonymous):

2 passes

jimthompson5910 (jim_thompson5910):

good

jimthompson5910 (jim_thompson5910):

so on average, you'd expect him to make two attempts til he gets a completion

OpenStudy (anonymous):

ok! c

OpenStudy (anonymous):

?

jimthompson5910 (jim_thompson5910):

one sec

OpenStudy (anonymous):

k

jimthompson5910 (jim_thompson5910):

looking for a calculator that will speed things up (because it's going to a pain to compute the CDF here)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

btw, which statements below are true? I had only III but im not sure: I. The expected value of a geometric random variable is determined by the formula np. II. If X is a geometric random variable and the probability of success is .85, then the probability distribution of X will be skewed left, since .85 is closer to 1 than 0. III. An important difference between binomial and geometric random variables is that there are a fixed number of trials in a binomial setting, and the number of trials varies in a geometric setting.

jimthompson5910 (jim_thompson5910):

hmm can't find one, we'll just have to do it this way

OpenStudy (anonymous):

ok

jimthompson5910 (jim_thompson5910):

when k = 1, P(X = k) = 0.54 when k = 2, P(X = k) = 0.2484 when k = 3, P(X = k) = 0.114264 when k = 4, P(X = k) = 0.05256144 when k = 5, P(X = k) = 0.0241782624 these probabilities add to 0.54+0.2484+0.114264+0.05256144+0.0241782624 = 0.9794037024 So basically P(X <= 5) = 0.9794037024 To find P(X > 5), just subtract it from 1 to get 1 - 0.9794037024 = 0.0205962976 So the probability that he has to throw more than 5 passes to get his first completion is 0.0205962976

OpenStudy (anonymous):

ok thanks i put P(X less than or equal to 5) = 0.9794037024 P(X>5) = 1 - 0.9794037024 = 0.0205963 <---

OpenStudy (anonymous):

can u tell me what the answer is to the prob i commented up there? i got III only, but im not sure..

jimthompson5910 (jim_thompson5910):

yeah that works, ok one sec

OpenStudy (anonymous):

k

jimthompson5910 (jim_thompson5910):

good, III looks like the only true statement

OpenStudy (anonymous):

kk thanks so much @jim_thompson5910 !!

jimthompson5910 (jim_thompson5910):

yw